# Brought a Gun to a Knife fight: Modern VFM Baselines Outgun Specialized Detectors on In-the-Wild AI Image Detection

### v1

@@ Abstract
While specialized detectors for AI-generated images excel on curated benchmarks, they fail catastrophically in real-world scenarios, as evidenced by their critically high false-negative rates on ‘in-the-wild’ benchmarks. Instead of crafting another specialized ‘knife’ for this problem, we bring a ‘gun’ to the fight: a simple linear classifier on a modern Vision Foundation Model (VFM). Trained on identical data, this baseline decisively ‘outguns’ bespoke detectors, boosting in-the-wild accuracy by a striking margin of over 20%.
Our analysis pinpoints the source of the VFM’s ‘firepower’: First, by probing text-image similarities, we find that recent VLMs (e.g., Perception Encoder, Meta CLIP2) have learned to align synthetic images with forgery-related concepts (e.g., ‘AI-generated’), unlike previous versions. Second, we speculate that this is due to data exposure, as both this alignment and overall accuracy plummet on a novel dataset scraped after the VFM’s pre-training cut-off date, ensuring it was unseen during pre-training. Our findings yield two critical conclusions: 1) For the real-world ‘gunfight’ of AI-generated image detection, the raw ‘firepower’ of an updated VFM is far more effective than the ‘craftsmanship’ of a static detector. 2) True generalization evaluation requires test data to be independent of the model’s entire training history, including pre-training.
@@ 1 Introduction
Fig. 1: Comparison of detection performance between modern VFMs and state-of-the-art forensics-specialized detectors on the GenImage and Chameleon.
@@
To investigate this issue, we conduct two analyses. First, focusing on the VLM-based backbones, we probe text-image similarities and find they have already aligned synthetic images with forgery-related concepts (e.g., ‘AI-generated’) well, a notable characteristic absent in earlier models. This provides a clear concept clue for their success. Second, to verify the broader hypothesis of data exposure for all tested VFMs, we collected a novel dataset of images generated after their pre-training cut-off date. As predicted, nearly all top models exhibit a noticeable performance drop. Thus, their remarkable performance appears to be a direct consequence of the model’s inadvertent training on the concept of contemporary AIGI within its pre-training data. This effectively equips them with a ’built-in’ effectiveness to in-the-wild content, allowing these general-purpose models to decisively outgun specialist detectors explicitly designed for this forensic task.
@@ 2.1 Experimental Setup
Together, these datasets provide a comprehensive and challenging testbed for evaluating a model’s true generalization capabilities.
@@
Method GenImage Chameleon Avg. R.Acc F.Acc Avg. CNNSpot (CVPR’20) [17] 64.2 99.7 0.08 56.9 FreqNet (AAAI’24) [18] 86.8 87.1 13.7 55.6 GramNet (CVPR’20) [19] 69.9 99.6 4.76 58.9 UnivFD (CVPR’23) [1] 88.8 97.8 3.18 57.2 NPR (ECCV’24) [20] 88.6 98.7 2.20 57.3 AIDE (ICLR’25) [5] 86.8 98.5 5.04 58.4 PPL (arxiv’25) [21] 97.2 68.1 64.7 66.6 OMAT (arxiv’25) [22] 94.6 90.2 33.9 66.1 NPLB (arxiv’25) [4] 97.1 79.4 65.5 73.4 CLIP (2021.2.26) 84.3 70.7 35.5 55.6 Siglip (2023.3.27) 88.6 70.8 49.3 61.6 Siglip2 (2025.2.21) 92.0 88.9 77.0 83.8 Meta CLIP (2023.9.28) 76.6 37.6 91.5 60.7 Meta CLIP-2 (2025.7.29) 89.5 84.8 87.9 91.8 PE (2025.4.17) 93.8 97.0 94.8 96.1 DINOv2 (2023.10.26) 85.3 62.9 58.1 60.8 DINOv3 (2025.8.14) 96.3 93.7 90.5 92.4
@@ 2.3 Validation on Additional In-the-Wild Datasets
Method WildRF Facebook Reddit Twitter Avg. NPR (ECCV’24) [20] 78.1 61.0 51.3 63.5 UnivFD (CVPR’23) [1] 49.1 60.2 56.5 55.3 FatFormer (CVPR’24) [23] 54.1 68.1 54.4 58.9 SAFE (KDD’25) [24] 50.9 74.1 37.5 57.2 C2P-CLIP (AAAI’25) [25] 54.4 68.4 55.9 59.6 AIDE (ICLR’25) [5] 57.8 71.5 45.8 58.4 DRCT (ICML’24) [2] 46.6 53.1 55.2 50.6 AlignedForensics (ICLR’25)[26] 89.4 69.1 81.8 80.1 DDA (arxiv’25) [27] 94.4 94.3 96.8 95.1 CLIP (2021.2.26) 77.2 72.3 68.9 72.8 Siglip (2023.3.27) 58.4 66.5 72.2 65.7 Siglip2 (2025.2.21) 79.9 84.7 89.6 84.7 Meta CLIP (2023.9.28) 63.8 74.1 77.9 71.9 Meta CLIP-2 (2025.7.29) 70.9 76.7 81.2 76.3 Pe (2025.4.17) 81.6 84.9 88.8 85.1 DINOv2 (2023.10.26) 71.3 71.0 71.3 71.2 DINOv3 (2025.8.14) 94.7 96.7 97.9 96.4
@@ 3 Analysis
3.1 Data Exposure Hypothesis: A Test on Post-Cutoff Data
Our primary speculation is that the superior in-the-wild performance of modern VFMs is a direct result of exposure to contemporary synthetic media during pre-training, rather than an intrinsic leap in generalization. To rigorously test this, we constructed a new dataset, WebAIG-25, designed to serve as a decisive blind test. We built it by collecting synthetic samples posted on the subreddit ‘r/midjourney‘ and real-world photographs from Unsplash. Crucially, all samples were uploaded by users between August 1, 2025, and September 10, 2025. This specific timeframe places the data after the pre-training data cutoff for most models under evaluation, making WebAIG-25 a powerful tool to disentangle true generalization from learned artifacts.
Table 4: Comparison results on WebAIG-25.
Method WebAIG-25 R.Acc F.Acc Avg. CLIP (2021.2.26) 52.1 71.1 61.6 Siglip (2023.3.27) 46.7 44.1 45.4 Siglip2 (2025.2.21) 84.8 55.8 70.3 Meta CLIP (2023.9.28) 59.7 55.6 57.7 Meta CLIP-2 (2025.7.29) 80.1 75.4 77.7 PE (2025.4.17) 100.0 91.7 95.9 DINOv2 (2023.10.26) 57.0 54.2 55.6 DINOv3 (2025.8.14) 98.6 91.4 95.0
The results on this ‘post-cutoff’ dataset, presented in Table 4, are revealing and offer strong support for our hypothesis. The most telling evidence comes from the modern VFM families. While the modern MetaCLIP-2 (77.7%) still outperforms its predecessor MetaCLIP (57.7%), its accuracy represents a stark drop from the 91.8% it achieved on Chameleon. A similar, significant drop is also observed for Siglip2. This pattern strongly suggests their high performance on older ‘in-the-wild’ sets is less an indicator of true generalization and more a result of prior data exposure. It is this very performance drop on verifiably unseen data that serves as the most compelling evidence, confirming that data exposure—not a fundamental leap in abstract generalization—is the dominant factor driving the remarkable success of modern VFMs on established benchmarks.
@@ 3.1 Data Exposure Hypothesis: A Test on Post-Cutoff Data
Method Chameleon SocialRF Top-1 Top-2 Top-1 Top-2 Matched Text Similarity Score Matched Text Similarity Score Matched Text Similarity Score Matched Text Similarity Score CLIP (2021.2.26) modern design 0.628 portrait 0.345 forged 0.332 urban 0.253 Siglip (2023.3.27) unaltered 0.318 original 0.212 edited 0.232 original 0.222 Siglip2 (2025.2.21) genuine 0.385 urban 0.250 portrait 0.202 vintage Meta CLIP (2023.9.28) AI generated 0.678 deepfake 0.091 AI generated 0.902 original 0.024 Meta CLIP-2 (2025.7.29) AI generated 0.828 deepfake 0.064 AI generated 0.924 deepfake 0.038 PE (2025.4.17) AI generated 0.861 deepfake 0.021 AI generated 0.943 deepfake 0.031 Method CommunityAI WebAIG-25 Top-1 Top-2 Top-1 Top-2 Matched Text Similarity Score Matched Text Similarity Score Matched Text Similarity Score Matched Text Similarity Score CLIP (2021.2.26) portrait 0.346 nature 0.326 urban 0.332 midjourney_images 0.260 Siglip (2023.3.27) real 0.281 AIGIBench 0.283 fake 0.278 original 0.246 Siglip2 (2025.2.21) urban 0.209 portrait 0.201 urban 0.212 portrait 0.194 Meta CLIP (2023.9.28) AI generated 0.726 deepfake 0.086 midjourney_images 0.604 AI generated 0.284 Meta CLIP-2 (2025.7.29) AI generated 0.858 deepfake 0.041 midjourney_images 0.621 AI generated 0.308 PE (2025.4.17) AI generated 0.878 CommunityAI 0.029 midjourney_images 0.722 AI generated 0.218
@@ 3.2 Probing VFM’s Latent Alignment with Forgery Concepts
To investigate the source of the superior in-the-wild performance observed in modern foundation models, we speculate that recent Vision-Language Models (VLMs) may have learned to associate synthetic images with forgery-related semantic concepts during pre-training. This would provide a powerful, high-level cue for detection that earlier models and non-language-aligned models like DINOv3 might lack.
To test this hypothesis, we designed a text-image similarity probing experiment exclusively for the VLM-based backbones. Specifically, we randomly sample images from several in-the-wild datasets. For each image, we compute its visual embedding using the VLM’s image encoder and measure its cosine similarity to a set of textual embeddings generated by the corresponding text encoder. These textual concepts are categorized into three groups:
  • Real and Fake-Related Keywords: [‘fake’, ‘real’, ‘AI generated’, ‘authentic’, ‘manipulated’, ‘synthetic’, etc],
  • Redundant Keywords: [‘sunset’, ‘landscape’, ‘portrait’, ‘abstract art’, ‘technology’, ‘nature’, etc],
  • Training Dataset Names: [‘GenImage’, ‘ADM’, ‘BigGAN’, ‘glide’, ‘Midjourney’, etc].
We randomly sampled 2,000 images from each in-the-wild dataset. Then calculated the similarity between these images and our pre-defined text pool, outputting the most textually similar terms for each image.
Table 5 reveals a significant generational divide in semantic understanding. Older models like CLIP and Siglip consistently fail to associate fake images with forgery-related concepts, instead matching them to generic terms like ‘portrait’ or ‘urban’. Their latent space essentially lacks a concept of "fakeness."
In stark contrast, modern VLMs like MetaCLIP-2 and Pe demonstrate a remarkably strong and consistent alignment. Across both established in-the-wild datasets and our new WebAIG-25, their top-matched concepts for fake images are overwhelmingly forgery-related terms like "AI generated" or "midjourney_images". This powerful semantic alignment, learned through exposure to contemporary web data, serves as a decisive shortcut for detection. It confirms that modern VLMs excel not just by recognizing visual artifacts, but by understanding the very concept of a "fake" image.
@@ 3.3 Visualization
The t-SNE visualizations in Fig. 2 offer a compelling qualitative explanation for the quantitative performance gaps in Tab. 1. On the curated GenImage dataset, both Meta CLIP-2 and CLIP exhibit similarly entangled feature spaces, which aligns with the comparable, high performance of both our VFM baselines (89.5%) and CLIP-based detectors like UnivFD (88.8%). However, this parity vanishes on the in-the-wild Chameleon dataset. There, the performance of UnivFD collapses to 57.2%, a failure visually mirrored in its CLIP backbone’s chaotic and inseparable feature space. In stark contrast, MetaCLIP2’s near-perfect accuracy (91.8%) is directly explained by its t-SNE plot, which reveals a remarkably clean separation between real and fake clusters. This direct link between feature separability and downstream performance confirms that MetaCLIP2’s advantage stems from a fundamentally more discriminative representation for in-the-wild content, acquired through pre-training exposure (Fig. 2).
@@ 4 Conclusion
This paper demonstrated that for in-the-wild AI-generated image detection, a simple linear classifier built upon a modern Vision Foundation Model (VFM) decisively outperforms specialized forensic detectors, akin to bringing a gun to a knife fight. We attribute this stark performance gap not to architectural novelty but to the VFM’s pre-training on vast, contemporary web data that includes synthetic media. Our findings have two critical implications: first, for practical forensics, leveraging the evolving representations of the latest foundation models is a more robust strategy than designing static detectors. Second, we call for a stricter evaluation protocol where test data is guaranteed to be novel to a model’s entire pre-training corpus to truly assess generalization.

### v2

@@ Abstract
While specialized detectors for AI-generated images excel on curated benchmarks, they fail catastrophically in real-world scenarios, as evidenced by their critically high false-negative rates on ‘in-the-wild’ benchmarks. Instead of crafting another specialized ‘knife’ for this problem, we bring a ‘gun’ to the fight: a simple linear classifier on a latest Vision Foundation Model (VFM). Trained on identical data, this baseline decisively ‘outguns’ bespoke detectors, boosting in-the-wild accuracy by a striking margin of over 20%.
Our analysis pinpoints the source of the VFM’s ‘firepower’: First, by probing text-image similarities, we find that recent VLMs (e.g., Perception Encoder, Meta CLIP-2) have learned to align synthetic images with forgery-related concepts (e.g., ‘AI-generated’), unlike their previous versions. Second, we speculate that this is due to data exposure, as performance plummets on a verifiably unseen dataset composed of post-cutoff synthetic images and private photographs. Our findings yield two critical insights: 1) For the real-world ‘gunfight’ of AI-generated image detection, the raw ‘firepower’ of an updated VFM is far more effective than the ‘craftsmanship’ of a static detector. 2) True generalization evaluation requires test data to be independent of the model’s entire training history, including pre-training.
@@ 1 Introduction
Fig. 1: Comparison of detection performance between modern VFMs and state-of-the-art forensics-specialized detectors on the GenImage[6] and Chameleon[5].
@@
To investigate this issue, we conduct two analyses. We first test all VFMs on a verifiably unseen dataset composed of post-cutoff synthetic media and private, personally captured photographs. As hypothesized, the performance of most modern VFMs drops significantly, revealing their reliance on prior data exposure. Building on this, our second analysis probes the underlying mechanism for Vision-Language Models. We find they have learned to align synthetic images with forgery-related concepts (e.g., ’AI-generated’), providing a semantic explanation for their effectiveness. This ’built-in’ knowledge, learned from web-scale data, allows these general-purpose models to decisively outgun specialist detectors explicitly designed for the forensic task.
@@ 2.1 Experimental Setup
Method GenImage Chameleon Avg. R.Acc F.Acc Avg. CNNSpot [17] 64.2 99.7 0.08 56.9 (-7.3) FreqNet [18] 86.8 87.1 13.7 55.6 (-31.2) GramNet [19] 69.9 99.6 4.76 58.9 (-11.0) UnivFD [1] 88.8 97.8 3.18 57.2 (-31.6) NPR [20] 88.6 98.7 2.20 57.3 (-31.3) AIDE [5] 86.8 98.5 5.04 58.4 (-28.4) PPL [21] 97.2 68.1 64.7 66.6 (-30.6) OMAT [22] 94.6 90.2 33.9 66.1 (-28.5) NPLB [4] 97.1 79.4 65.5 73.4 (-23.7) CLIP (2021.2.26) 84.3 70.7 35.5 55.6 (-28.7) Siglip (2023.3.27) 88.6 70.8 49.3 61.6 (-27.0) Siglip2 (2025.2.21) 92.0 88.9 77.0 83.8 (-8.2) MetaCLIP (2023.9.28) 76.6 37.6 91.5 60.7 (-15.9) MetaCLIP-2 (2025.7.29) 89.5 84.8 87.9 91.8 (+2.3) PE (2025.4.17) 93.8 97.0 94.8 96.1 (+2.3) DINOv2 (2023.10.26) 85.3 62.9 58.1 60.8 (-24.5) DINOv3 (2025.8.14) 96.3 93.7 90.5 92.4 (-3.9)
@@ 2.3 Validation on Additional In-the-Wild Datasets
Method Facebook Reddit Twitter Avg. NPR (ECCV’24) [20] 78.1 61.0 51.3 63.5 UnivFD (CVPR’23) [1] 49.1 60.2 56.5 55.3 FatFormer (CVPR’24) [23] 54.1 68.1 54.4 58.9 SAFE (KDD’25) [24] 50.9 74.1 37.5 57.2 C2P-CLIP (AAAI’25) [25] 54.4 68.4 55.9 59.6 AIDE (ICLR’25) [5] 57.8 71.5 45.8 58.4 DRCT (ICML’24) [2] 46.6 53.1 55.2 50.6 AlignedForensics (ICLR’25)[26] 89.4 69.1 81.8 80.1 CLIP (2021.2.26) 77.2 72.3 68.9 72.8 Siglip (2023.3.27) 58.4 66.5 72.2 65.7 Siglip2 (2025.2.21) 79.9 84.7 89.6 84.7 Meta CLIP (2023.9.28) 63.8 74.1 77.9 71.9 Meta CLIP-2 (2025.7.29) 70.9 76.7 81.2 76.3 PE (2025.4.17) 81.6 84.9 88.8 85.1 DINOv2 (2023.10.26) 71.3 71.0 71.3 71.2 DINOv3 (2025.8.14) 94.7 96.7 97.9 96.4
@@ 3 Analysis
3.1 Testing for Data Exposure with Verifiably Unseen Data
Our primary speculation is that the superior in-the-wild performance of modern VFMs is a direct result of exposure to contemporary synthetic media during pre-training, rather than an intrinsic leap in generalization. To rigorously test this, we constructed a new dataset designed to be a decisive blind test, where both real and fake images are verifiably unseen. Its synthetic and real portions are both designed to be verifiably unseen. The former comprises r/midjourney images posted after all model training cut-off date, while the latter consists of private, personally captured photographs of campus scenes, guaranteeing the entire dataset’s absence from any pre-training corpus. This setup allows for a rigorous evaluation of true generalization, free from the confounding variable of data overlap between pre-training corpora and public benchmarks.
Table 4: Comparison results on WebAIG-25 dataset.
Method R.Acc F.Acc Avg. CNNSpot (CVPR’20) [17] 97.1 8.3 52.7 UnivFD (CVPR’23) [1] 59.2 29.8 44.5 NPR (ECCV’24) [20] 93.5 34.9 64.2 SAFE (KDD’25) [24] 99.7 21.3 60.5 FreqNet (AAAI’24) [18] 67.0 26.0 46.5 C2P-CLIP (AAAI’25) [25] 84.7 31.0 57.9 PatchShuffle (NeurIPS’24) [27] 54.9 45.0 50.0 CLIP (2021.2.26) 46.7 71.1 58.9 Siglip (2023.3.27) 31.3 44.1 37.7 Siglip2 (2025.2.21) 69.5 55.8 62.7 Meta CLIP (2023.9.28) 36.2 55.6 45.9 Meta CLIP-2 (2025.7.29) 27.4 75.4 51.4 PE (2025.4.17) 44.4 91.7 68.1 DINOv2 (2023.10.26) 40.9 54.2 47.6 DINOv3 (2025.8.14) 97.5 91.4 94.5
The results on this verifiably unseen dataset (Table 4) are striking, revealing distinct and opposing biases rooted in different training paradigms. Specialized detectors show a strong bias towards "real", achieving high accuracy on our private real photos while catastrophically failing on novel fakes. Conversely, modern Vision-Language Models, trained via text-image contrastive learning, exhibit an opposite bias: they excel at identifying novel fakes but struggle with our out-of-distribution real photos, suggesting their core training objective of aligning images with text descriptions limits their ability to generalize to visual distributions, like our private photos, that fall outside this vast but specific semantic space. Amidst these widespread biases, the self-supervised DINOv3 emerges as the sole exception, achieving exceptional performance on both real and fake images. This suggests that while data exposure dictates the biases of most models, only DINOv3’s robust, fine-grained visual representations, learned without textual shortcuts, can generalize to both unseen real and synthetic distributions.
@@ 3.1 Testing for Data Exposure with Verifiably Unseen Data
Method Chameleon SocialRF Top-1 Top-2 Top-1 Top-2 Matched Text Similarity Score Matched Text Similarity Score Matched Text Similarity Score Matched Text Similarity Score CLIP (2021.2.26) modern design 0.628 portrait 0.345 forged 0.332 urban 0.253 Siglip (2023.3.27) unaltered 0.318 original 0.212 edited 0.232 original 0.222 Siglip2 (2025.2.21) genuine 0.385 urban 0.250 portrait 0.202 vintage 0.191 Meta CLIP (2023.9.28) AI generated 0.678 deepfake 0.091 AI generated 0.902 original 0.024 Meta CLIP-2 (2025.7.29) AI generated 0.828 deepfake 0.064 AI generated 0.924 deepfake 0.038 PE (2025.4.17) AI generated 0.861 deepfake 0.021 AI generated 0.943 deepfake 0.031 Method CommunityAI WebAIG-25 Top-1 Top-2 Top-1 Top-2 Matched Text Similarity Score Matched Text Similarity Score Matched Text Similarity Score Matched Text Similarity Score CLIP (2021.2.26) portrait 0.346 nature 0.326 urban 0.332 midjourney_images 0.260 Siglip (2023.3.27) real 0.281 AIGIBench 0.283 fake 0.278 original 0.246 Siglip2 (2025.2.21) urban 0.209 portrait 0.201 urban 0.212 portrait 0.194 Meta CLIP (2023.9.28) AI generated 0.726 deepfake 0.086 midjourney_images 0.604 AI generated 0.284 Meta CLIP-2 (2025.7.29) AI generated 0.858 deepfake 0.041 midjourney_images 0.621 AI generated 0.308 PE (2025.4.17) AI generated 0.878 CommunityAI 0.029 midjourney_images 0.722 AI generated 0.218
@@ 3.2 Probing VFM’s Latent Alignment with Forgery Concepts
To investigate the mechanism behind the success of Vision-Language Models (VLMs), we hypothesize that their pre-training exposure has allowed them to learn a powerful semantic shortcut: associating synthetic images with high-level, forgery-related textual concepts. This provides a potential high-level cue for detection that we can directly probe.
To test this specific hypothesis, we designed a text-image similarity probing experiment applicable only to VLM-based backbones, as it relies on their inherent text-encoder capabilities. This analysis is therefore not applicable to self-supervised models like DINOv3, whose remarkable performance must stem from a different, purely visual mechanism. First, we constructed a comprehensive text pool categorized into three groups:
  • Forgery-Related Concepts: [‘fake’, ‘real’, ‘AI generated’, ‘authentic’, ‘manipulated’, ‘synthetic’, etc.]
  • Content-Related Concepts: [‘sunset’, ‘landscape’, ‘portrait’, ‘abstract art’, ‘technology’, ‘nature’, etc.]
  • Source-Related Concepts: [‘GenImage’, ‘ADM’, ‘BigGAN’, ‘glide’, ‘Midjourney’, etc.]
For the experiment, we randomly sampled 2,000 fake images from each in-the-wild dataset. We then computed the cosine similarity between each image’s visual embedding and the textual embeddings of all concepts in our pool. The top-matching text concepts for each VLM are reported to reveal their latent semantic alignments.
Table 5 reveals a significant generational divide in semantic understanding. Older models like CLIP and Siglip consistently fail to associate fake images with forgery-related concepts, instead matching them to generic terms like ‘portrait’ or ‘urban’. Their latent space essentially lacks a concept of "fakeness." In stark contrast, modern VLMs like Meta CLIP-2 and PE demonstrate a remarkably strong and consistent alignment. Across both established in-the-wild datasets and our new WebAIG-25, their top-matched concepts for fake images are overwhelmingly forgery-related terms like "AI generated" or "midjourney_images". It confirms that modern VLMs excel not just by recognizing visual artifacts, but by understanding the very concept of a "fake" image.
@@ 3.3 Visualization
The t-SNE visualizations in Fig. 2 offer a compelling qualitative explanation for the quantitative performance gaps in Tab. 1. On the curated GenImage dataset, both Meta CLIP-2 and CLIP exhibit similarly entangled feature spaces, which aligns with the comparable, high performance of both our VFM baselines (89.5%) and CLIP-based detectors like UnivFD (88.8%). However, this parity vanishes on the in-the-wild Chameleon dataset. There, the performance of UnivFD collapses to 57.2%, a failure visually mirrored in its CLIP backbone’s chaotic and inseparable feature space. In stark contrast, Meta CLIP-2’s near-perfect accuracy (91.8%) is directly explained by its t-SNE plot, which reveals a remarkably clean separation between real and fake clusters. This direct link between feature separability and downstream performance confirms that Meta CLIP-2’s advantage stems from a fundamentally more discriminative representation for in-the-wild content, acquired through pre-training exposure (Fig. 2).
@@ 4 Conclusion
This paper demonstrated that for in-the-wild AI-generated image detection, a simple linear classifier built upon a modern Vision Foundation Model (VFM) decisively outperforms specialized forensic detectors, akin to bringing a gun to a knife fight. We attribute this stark performance gap not to architectural novelty but to the VFM’s pre-training on vast, contemporary web data that includes synthetic media. Our findings have two critical implications: First, for practical forensics, leveraging the evolving representations of the latest foundation models is a more robust strategy than designing static detectors. Second, for academic research, a stricter evaluation protocol that uses test data guaranteed to be novel to a model’s entire pre-training corpus is paramount to truly assess generalization.